We wish to compare the prevalence of parasites in Mediterranean fish to the prevalence of parasites in Atlantic fish. In the Mediterranean Sea 500 brill were captured and dissected; 200 were found to contain parasites   In the Atlantic 200 brill were captured and dissected; 40 were found to have parasites. In what follows, let p1 denotes the infestation rate for the Mediterranean brill and p2 denote the infestation rate for the Atlantic brill. Find the point estimate for the difference p1   - p2.

0.35

160

0.20

0.10

0.60
Answer #1

Solution:

p1 = x / n = 200 / 500

p1 = 0.40

p2 = x / n = 40 /200

p2 = 0.20

Point estimate = p1 - p2 = 0.40 - 0.20

= 0.20